OpenStudy (anonymous):

Solve using the substitution method. Show your work. If the system has no solution or an infinite number of solutions, state this. 4a + 20b = 8 8a = -40b + 16

OpenStudy (anonymous):

Substitution means that you use an expression of one variable and substitute it into the other equation... sort of like elimination, this allows you to solve for one variable first, then the other.

OpenStudy (lukebluefive):

Divide both sides of the first equation by 4: 4a + 20b = 8 a + 5b = 2 Subtract 5b from both sides a = 2 - 5b Now take the second equation and divide both sides by 8: a = -5b + 2 Since these two equations are identical, there are an infinite number of solutions.

OpenStudy (anonymous):

no body wants to help lol

OpenStudy (lukebluefive):

There are an infinite number of solutions since the two lines are on top of each other, and since lines go on indefinitely, there are infinitely many points where they intersect.

OpenStudy (anonymous):

@LukeBlueFive yours is totally correct, but I think the problem requires substitution...

OpenStudy (anonymous):

@JakeV8 so do i solve for a first?

OpenStudy (lukebluefive):

Well, you can substitute a = 2 - 5b into 8a = -40b + 16: 8(2 - 5b) = -40b + 16 16 - 40b = -40b + 16 Then, infinitely many solutions for the reasons I stated above.

OpenStudy (anonymous):

@JakeV8 and @LukeBlueFive sorry i was frozen =(

OpenStudy (lukebluefive):

@jdorta1 Same here.

OpenStudy (anonymous):

that was weird